OpenStudy (anonymous):

The formula r=2sqrt(5L) can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length L, in feet. How far will a car skid at 20 mph?

OpenStudy (anonymous):

\[20 = 2 \sqrt{5L}\]\[10 = \sqrt {5L}\]\[10^{2} = 5L\]\[L = 100/5 = 20\]

OpenStudy (anonymous):

So how many feet is it? 100/5?

OpenStudy (anonymous):

transform 20 mi into feet 20 mi = 105 600 feet

OpenStudy (anonymous):

okay, what do i do next?

OpenStudy (anonymous):

that's yur answer..

OpenStudy (anonymous):

105 600 feet is my answer?

OpenStudy (anonymous):

a car that has left skid marks of length L, in feet. How far will a car skid at 20 mph? -> 20 mi = 105 600 feet

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

what about at 80mph